im trying to find the equation of the line containing the points (0,8) and 6x+7y=4
Answer by mananth(16946)

6x+7y=4
write in slope intercept form
7y=-6x+4
/7
y=(-6/7)x+(4/7)
compare with y = mx+b
slope m = (-6/7)
the line passes through (0,8)
the general equation when slope and point is given
(y-y1)=m(x-x1) where (x1,y1) is the point.
(x1,y1) = (0,8)
(y-8)=(-6/7)(x-0)
y-8=-6/7 x
y+(6/7)x=8
multiply by 7
7y+6x= 56 is the equation of the line